# RouteKG: A knowledge graph-based framework for route prediction on road networks

### v2

@@ Abstract
Short-term route prediction on road networks allows us to anticipate the future trajectories of road users, enabling a plethora of intelligent transportation applications such as dynamic traffic control or personalized route recommendation. Despite recent advances in this area, existing methods focus primarily on learning sequential transition patterns, neglecting the inherent spatial structural relations in road networks that can affect human routing decisions. To fill this gap, this paper introduces RouteKG, a novel Knowledge Graph-based framework for route prediction. Specifically, we construct a Knowledge Graph on the road network, thereby learning and leveraging spatial relations, especially moving directions, which are crucial for human navigation. Moreover, an $n$-ary tree-based algorithm is introduced to efficiently generate top-$K$ routes in a batch mode, enhancing scalability and computational efficiency. To further optimize the prediction performance, a rank refinement module is incorporated to fine-tune the candidate route rankings. The model performance is evaluated using two real-world vehicle trajectory datasets from two Chinese cities, Chengdu and Shanghai, under various practical scenarios. The results demonstrate a significant improvement in accuracy over baseline methods.We further validate our model through a case study that utilizes the pre-trained model as a simulator for real-time traffic flow estimation at the link level. The proposed RouteKG promises wide-ranging applications in vehicle navigation, traffic management, and other intelligent transportation tasks.
keywords:
@@
Route prediction, Knowledge graph, Road network representation, Trajectory data mining, Geospatial AI
1 Introduction
@@
In intelligent transportation systems (ITS), with the increasing prevalence of mobile sensors (e.g., GPS devices) and vehicular communication technologies, the ability to predict road users’ future routes is not merely a convenience but a necessity to support a range of applications such as vehicle navigation (Ziebart et al. 2008), traffic management (Li et al. 2020) and location-based recommendation (Kong et al. 2017, Tang et al. 2022). There are generally two types of route prediction tasks. On the one hand, for transport planning applications, it is often required to predict the complete route (as a sequence of road links) from the origin to destination. This is typically referred to as route choice modeling in the literature (Prato 2009), where the destination information has to be given. On the other hand, for real-time ITS applications, the destination information may not be available, and it is usually adequate to predict the near-future route trajectory of a moving agent based on the observed trajectory so far. This study focuses on the latter, which we call the short-term route prediction problem.
@@
Numerous methods have been proposed in the literature to tackle this problem. Recent works typically use Recurrent Neural Networks (RNNs) (Rumelhart et al. 1986), especially Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber 1997) and Gated Recurrent Unit (GRU) (Cho et al. 2014), to capture sequential dependencies in trajectory data (Alahi et al. 2016, Mo et al. 2023). Most existing models focus primarily on learning sequential patterns for route prediction, often overlooking the inherent spatial structure in road networks that can affect human routing decisions. To address this issue, some studies have started to leverage Graph Neural Networks (GNNs) to encode road networks for improved prediction of vehicle trajectories (Liang & Zhao 2021) and traffic conditions (Zhao et al. 2019, Li et al. 2017). However, these methods still treat road networks merely as generic graphs, oversimplifying their structure and disregarding crucial geographical attributes and spatial factors.
@@
As a type of spatial network, road networks consist of a set of spatial entities (e.g., intersections, links, etc.) organized in a way to facilitate traffic flows in a mostly 2-dimensional space. The relationships between these entities can be described by a set of spatial factors such as direction, distance, and connectivity. For example, one of the important spatial factors to consider in routing problems is the direction of travel (i.e., goal direction). It has been widely recognized in the navigation and cognitive psychology literature that humans utilize directional cues to navigate their environment (Etienne & Jeffery 2004, Chrastil & Warren 2015). Existing short-term route prediction models, however, often ignore the directional factor or incorporate it as a side feature, thus possibly leading to sub-optimal model performance. These limitations highlight the need for a more spatially explicit model to learn and incorporate these spatial relations throughout road networks.
@@
There are other challenges for short-term route prediction on road networks. Firstly, most existing methods focus on generating a single predicted route (Rathore et al. 2019, Yan et al. 2022). However, due to the inherent uncertainties, providing multiple route predictions can have more practical implications. For instance, traffic managers can optimize real-time traffic flow by considering multiple potential routes of moving vehicles, and transportation system users can benefit from having a wider variety of routing options. Secondly, as road networks grow in size, scalability becomes a challenge for GNN-based methods (Hamilton et al. 2017), as they require substantial computational and memory resources, limiting their applicability to large-scale networks. Lastly, the prediction performance is heavily dependent on the availability of information about the destination (or goal). Generally, potential performance enhancements can be achieved by incorporating goal information into route prediction models. The availability of goal information can vary, including (1) no information, (2) goal direction only, and (3) complete goal information. These varying degrees of goal information availability can impact the route prediction to different extents, but no existing studies have carried out a comprehensive evaluation across all these scenarios.
@@
With the aforementioned challenges, we propose a novel model, “RouteKG”, which leverages the potential of Knowledge Graphs (KGs) (Wang et al. 2014) to encode road networks for short-term route prediction. Unlike existing models that rely on sequence-to-sequence (seq2seq) structures (Sutskever et al. 2014), our approach interprets route prediction as a Knowledge Graph Completion (KGC) task (Chen et al. 2020). Specifically, we propose a Knowledge Graph Module that can predict the future links (tail entities) a user might traverse based on current links (head entities) and moving directions (relations) without solely relying on sequential structures. The module explicitly incorporates the goal moving direction (estimated or actual) into the future route prediction process, better aligning with the intrinsic nature of human navigation. In addition, we employ a Route Generation Module to efficiently generate top-$K$ route candidates, and a Rank Refinement Module that can model the dependencies between different links within each predicted route to rerank the route candidates for their consistency, resulting in the final top-$K$ predictions. Our proposed KG-based framework can effectively model the spatial relations, thus outperforming existing baselines by a large margin, and could benefiting a range of other transportation or routing tasks. To summarize, this paper contributes to the literature as follows:
@@
  • We introduce RouteKG, a novel KG-based modeling framework for short-term route prediction. In this approach, we adapt the KG to represent road networks and reformulate the route prediction problem as a KGC task. Therefore, we can leverage the road network and route representations learned from the KG to enhance prediction accuracy and interpretability.
@@
2 Literature Review
2.1 Trajectory Prediction
2.1.1 Motion Prediction
Motion prediction, which anticipates an agent’s future trajectory from past movements, is central to autonomous driving systems (Yurtsever et al. 2020, Lefèvre et al. 2014). Its importance has amplified with advancements in autonomous driving and robot navigation, improving safety and efficiency by mitigating collision risks and boosting performance (Rudenko et al. 2020). However, the dynamic and uncertain nature of agents’ movements presents unique challenges (Paravarzar & Mohammad 2020). Motion prediction methods can generally be divided into two broad categories: classic and deep learning-based, each with unique advantages and limitations.
Classic methods leverage mathematical models grounded in physics and geometry to focus on the deterministic aspects of an agent’s motion, offering simplicity, interpretability, and efficiency (Helbing & Molnar 1995). Yet, these methods struggle to capture the stochastic behavior of agents in complex environments (Huang et al. 2022).
On the other hand, deep learning-based motion prediction methods leverage neural networks’ power to model the complexities of agent behavior (Alahi et al. 2016). These methods aim to learn the intricate, often non-linear, relationships between different influencing factors from large-scale data. Approaches such as RNNs and Generative Adversarial Networks (GANs) are commonly used (Gupta et al. 2018, Sadeghian et al. 2019, Gu et al. 2021). Recent efforts employ diffusion process (Ho et al. 2020) simulate the process of human motion variation from indeterminate to determinate (Gu et al. 2022). The advantage of deep learning methods is their ability to capture the underlying patterns and subtleties that traditional mathematical models might miss. However, they require extensive computational resources and large amounts of training data, and often lack the interpretability of classic methods (Rudenko et al. 2020).
2.1.2 Route Prediction
@@
Traditional methods utilize shortest path-based methods such as the Dijkstra’s algorithm (Dijkstra 1959), Bellman-Ford, and A* (Hart et al. 1968) for route prediction tasks. However, these dynamic programming-based methods require destination information to generate potential routes for trajectory prediction. As the destination information is often unavailable for short-term route prediction, other works have employed Kalman Filters (Abbas et al. 2020) or Hidden Markov Models (HMMs) (Simmons et al. 2006, Ye et al. 2016) to predict users’ destinations and routes. Nevertheless, these methods struggle to model long-term temporal dependencies due to relatively simple model structures.
In comparison, deep learning-based methods have outperformed traditional methods in prediction tasks, exhibiting superior ability in modeling spatial-temporal dependencies. The RNN-based encoder-decoder trajectory representation learning framework (Fu & Lee 2020) can adapt to tasks such as trajectory similarity measurement, travel time prediction, and destination prediction. Other studies have utilized Graph Convolutional Networks (GCN) and attention mechanisms to refine trajectory representation for prediction purposes (Shao et al. 2021). Some studies have proposed models for tasks ranging from predicting the next link using historical trajectories (Liu et al. 2022) to enhancing route prediction through pre-training and contrastive learning (Yan et al. 2022). Furthermore, some models are designed for road network-constrained trajectory recovery, capable of recovering fine-grained points from low-sampling records (Ren et al. 2021, Chen, Zhang, Sun & Zheng 2022).
@@
2.2 Knowledge Graph
2.2.1 Knowledge Graph Completion
The rapidly expanding interest in KGs has fueled the advancement in tasks like recommender systems, question answering, and semantic search, given their ability to provide structured and machine-interpretable knowledge about real-world entities and their relations (Noy et al. 2019, Sheth et al. 2019, Paulheim 2017). Despite their immense potential, a critical problem is the inherent incompleteness of information, making KGC an important and burgeoning research area. KGC refers to inferring missing or incomplete information in a KG by predicting new relationships between entities based on existing information (Chen et al. 2020).
Earlier studies on KGC typically employed statistical relational learning (SRL) methods, such as Markov Logic Networks (MLN) (Richardson & Domingos 2006) and Probabilistic Soft Logic (PSL) (Bach et al. 2017). These methods demonstrate effectiveness in capturing complex dependencies but need to improve scalability due to the need to specify all possible rules manually. More scalable machine learning approaches, especially those involving embeddings, have been proposed to overcome these limitations in recent years. TransE is a seminal model in this line, which models relations as translations in the entity embedding space (Bordes et al. 2013). Follow-up models such as TransH (Wang et al. 2014), TransR (Lin et al. 2015), and TransD (Ji et al. 2015) were subsequently proposed to handle complex relational data by introducing hyperplanes, relation-specific spaces, or dynamic mapping matrices respectively. Meanwhile, tensor factorization-based models like RESCAL (Nickel et al. 2011), DistMult (Yang et al. 2014), and ComplEx (Trouillon et al. 2016) have been developed, aiming to capture the complex correlations between entities and relations. These models generally perform well but can be computationally intensive. More recently, models based on GNNs have shown promising results for KGC. Models such as R-GCN (Schlichtkrull et al. 2018) and CompGCN (Vashishth et al. 2019) have achieved competitive results by modeling KGs as multi-relational graphs and learning from both the graph structure and node attributes.
@@ 2.2.1 Knowledge Graph Completion
2.2.2 Mobility Knowledge Graph
@@
Tan et al. (2021) devised a KG for urban traffic systems to uncover the implicit relationships amongst traffic entities and thereby unearth valuable traffic knowledge. Similarly, Zhuang et al. (2017) constructed an urban movement KG using GPS trajectory data and affirmed the practicality of their model by predicting the level of user attention directed towards various city locations. Zhao et al. (2020) put forth a generalized framework for multi-source spatiotemporal data analysis, underpinned by KG embedding, intending to discern the network structure and semantic relationships embedded within multi-source spatiotemporal data. Several studies have focused on building KGs grounded on geographical information and human mobilities for various applications, such as predicting subsequent locations (i.e., Point of Interest recommendation) (Liu et al. 2021, Rao et al. 2022, Wang et al. 2021), modeling event streams (Wang et al. 2020), learning user similarity (Zhang et al. 2023), forecasting destinations (Li et al. 2022, Chi et al. 2022), and performing epidemic contact tracing (Chen, Zhang, Qian & Li 2022).
@@
3 Preliminaries
In this section, we introduce definitions and the problem formulation in Section 3.1. All the notations used in this paper are listed in the A.
3.1 Problem Formulation
@@
A map-matched route $x$ of length $m$ is a sequence of links, $x=\left\{e^{1},e^{2},...,e^{m}\right\},x\in\mathcal{X}$. For every consecutive pair of links $(e^{i},e^{i+1})$, there exists a node $v$ in the graph $\mathbf{G}$ that connects the two edges. The set of all map-matched routes is denoted as $\mathcal{X}$. The $i$-th route can be partitioned into an observed route $x^{o}_{i}=\{e^{j}_{i}\}^{\Gamma}_{j=1}$ with length $\Gamma$ and a future route $x^{f}_{i}=\{e^{j}_{i}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$ with length Γsuperscriptnormal-$\Gamma^{\prime}$, where $\mathcal{X}^{o}=\{x_{i}^{o}\}_{i=1}^{|\mathcal{X}|}$ and $\mathcal{X}^{f}=\{x_{i}^{f}\}_{i=1}^{|\mathcal{X}|}$.
Given above definitions, the short-term route prediction (or route prediction for short) problem can be broadly defined as the task of predicting the future route based on observed routes. However, as discussed in Section 1, the availability of goal information plays a pivotal role in routing tasks. In some scenarios, no goal information is available. In other cases, we may know the rough direction of the destination, or its exact location. The degree of goal information inclusion can greatly influence the specific formulation of route prediction Dendorfer et al. (2020). Remarkably, no existing work has undertaken an exhaustive evaluation encompassing all these distinct scenarios. Consequently, in this study, we categorize the route prediction problem into three subproblems:

### v3

@@ Abstract
Short-term route prediction on road networks allows us to anticipate the future trajectories of road users, enabling various applications ranging from dynamic traffic control to personalized navigation. Despite recent advances in this area, existing methods focus primarily on learning sequential transition patterns, neglecting the inherent spatial relations in road networks that can affect human routing decisions. To fill this gap, this paper introduces RouteKG, a novel Knowledge Graph-based framework for route prediction. Specifically, we construct a Knowledge Graph on the road network to encode spatial relations, especially moving directions that are crucial for human navigation. Moreover, an n-ary tree-based algorithm is introduced to efficiently generate top-K routes in a batch mode, enhancing computational efficiency. To further optimize the prediction performance, a rank refinement module is incorporated to fine-tune the candidate route rankings. The model performance is evaluated using two real-world vehicle trajectory datasets from two Chinese cities under various practical scenarios. The results demonstrate a significant improvement in accuracy over baseline methods. We further validate the proposed method by utilizing the pre-trained model as a simulator for real-time traffic flow estimation at the link level. RouteKG holds great potential for transforming vehicle navigation, traffic management, and a variety of intelligent transportation tasks, playing a crucial role in advancing the core foundation of intelligent and connected urban systems.
Index Terms:
@@
Route prediction, Road network representation, Knowledge graph, Intelligent transportation systems
I Introduction
@@
In intelligent transportation and urban systems, with the increasing prevalence of mobile sensors (e.g., GPS devices) and vehicular communication technologies, the ability to predict road users’ future routes is not merely a convenience but a necessity to support a range of applications such as vehicle navigation [1], traffic management [2], accident prediction [3] and location-based recommendation [4, 5, 6]. There are generally two types of route prediction tasks. On the one hand, for transport planning applications, it is often required to predict the complete route (as a sequence of road links) from the origin to the destination. This is typically referred to as route choice modeling in the literature [7], where the destination information has to be given. On the other hand, for real-time ITS applications, the destination information may not be available, and it is usually adequate to predict the near-future route trajectory of a moving agent based on the observed trajectory so far. This study focuses on the latter, which we call the short-term route prediction problem.
@@
Numerous methods have been proposed in the literature to tackle this problem. Recent works typically use Recurrent Neural Networks (RNNs) [8], especially Long Short-Term Memory (LSTM) [9] and Gated Recurrent Unit (GRU) [10], to capture sequential dependencies in trajectory data [11, 12]. Most existing models focus primarily on learning sequential patterns for route prediction, often overlooking the inherent spatial structure in road networks that can affect human routing decisions. To address this issue, some studies have started to leverage Graph Neural Networks (GNNs) [13] to encode road networks for improved prediction of vehicle trajectories [14] and traffic conditions [15, 16, 17]. However, these methods still treat road networks merely as generic graphs, oversimplifying their structure and disregarding crucial geographical attributes and spatial factors.
@@
As a type of spatial network, road networks consist of a set of spatial entities (e.g., intersections, links, etc.) organized in a way to facilitate traffic flows in a mostly 2-dimensional space. The relationships between these entities can be described by a set of spatial factors such as direction, distance, and connectivity. For example, one of the important spatial factors to consider in routing problems is the direction of travel (i.e., goal direction). It has been widely recognized in the navigation and cognitive psychology literature that humans utilize directional cues to navigate their environment [18, 19]. Existing short-term route prediction models, however, often neglect the inherent static spatial structural relationships and the dynamic relations between these networks and human-intended moving directions, thus possibly leading to sub-optimal model performance. These limitations highlight the need for a more spatially explicit model to accurately learn and incorporate spatial relationships and human-intended moving directions within road networks.
@@
There are other challenges for short-term route prediction on road networks. Firstly, most existing methods focus on generating a single predicted route [20, 21]. However, due to the inherent uncertainties, providing multiple route predictions can have more practical implications. For instance, traffic managers can optimize real-time traffic flow by considering multiple potential routes of moving vehicles, and transportation system users can benefit from having a wider variety of routing options. Secondly, as road networks grow in size, scalability becomes a challenge for GNN-based methods [22], as they require substantial computational and memory resources, limiting their applicability to large-scale networks. Lastly, the prediction performance is heavily dependent on the availability of information about the destination (or goal). Generally, potential performance enhancements can be achieved by incorporating goal information into route prediction models. The availability of goal information can vary, including (1) no information, (2) goal direction only, and (3) complete goal information. These varying degrees of goal information availability can impact the route prediction to different extents, but no existing studies have carried out a comprehensive evaluation across all these scenarios.
@@
With the aforementioned challenges, we propose a novel model, “RouteKG”, which leverages the potential of Knowledge Graphs (KGs) [23] to encode road networks for short-term route prediction. Unlike existing models that rely on sequence-to-sequence (seq2seq) structures [24], our approach interprets route prediction as a Knowledge Graph Completion (KGC) task [25]. Specifically, we propose a Knowledge Graph Module that can predict the future links (tail entities) a user might traverse based on current links (head entities) and moving directions (relations) without solely relying on sequential structures. The module explicitly incorporates the goal moving direction (estimated or actual) into the future route prediction process, better aligning with the intrinsic nature of human navigation. In addition, we employ a Route Generation Module to efficiently generate top-$K$ route candidates, and a Rank Refinement Module that can model the dependencies between different links within each predicted route to rerank the route candidates for their consistency, resulting in the final top-$K$ predictions. Our proposed KG-based framework can effectively model the spatial relations, thus outperforming existing baselines by a large margin, and could benefiting a range of other transportation or routing tasks. To summarize, this paper contributes to the literature as follows:
@@
  • We introduce RouteKG, a novel KG-based modeling framework for short-term route prediction. In this approach, we adapt the KG to represent road networks and reformulate the route prediction problem as a KGC task. Then, the learned road network and route representations can enhance prediction accuracy and interpretability.
@@
II Literature Review
II-A Trajectory Prediction
II-A1 Motion Prediction
Motion prediction, which anticipates an agent’s future trajectory from past movements, is central to autonomous driving systems [26, 27]. Its importance has amplified with advancements in autonomous driving and robot navigation, improving safety and efficiency by mitigating collision risks and boosting performance [28]. However, the dynamic and uncertain nature of agents’ movements presents unique challenges [29]. Motion prediction methods can generally be divided into two broad categories: classic and deep learning-based, each with unique advantages and limitations.
Classic methods leverage mathematical models grounded in physics and geometry to focus on the deterministic aspects of an agent’s motion, offering simplicity, interpretability, and efficiency [30]. Yet, these methods struggle to capture the stochastic behavior of agents in complex environments [31].
On the other hand, deep learning-based motion prediction methods leverage neural networks’ power to model the complexities of agent behavior [11]. These methods aim to learn the intricate, often non-linear, relationships between different influencing factors from large-scale data. Approaches such as RNNs and Generative Adversarial Networks (GANs) are commonly used [32, 33, 34]. Recent efforts employ diffusion process [35] simulate the process of human motion variation from indeterminate to determinate [36]. The advantage of deep learning methods is their ability to capture the underlying patterns and subtleties that traditional mathematical models might miss. However, they require extensive computational resources and large amounts of training data, and often lack the interpretability of classic methods [28].
II-A2 Route Prediction
@@
Traditional methods utilize shortest path-based methods such as Dijkstra’s algorithm [37], Bellman-Ford, and A* [38] for route prediction tasks. However, these dynamic programming-based methods require destination information to generate potential routes for trajectory prediction. As the destination information is often unavailable for short-term route prediction, other works have employed Kalman Filters [39] or Hidden Markov Models (HMMs) [40, 41] to predict users’ destinations and routes. Nevertheless, these methods struggle to model long-term temporal dependencies due to relatively simple model structures.
In comparison, deep learning-based methods have outperformed traditional methods in prediction tasks, exhibiting superior ability in modeling spatial-temporal dependencies. The RNN-based encoder-decoder trajectory representation learning framework [42] can adapt to tasks such as trajectory similarity measurement, travel time prediction, and destination prediction. Other studies have utilized Graph Convolutional Networks (GCN) and attention mechanisms to refine trajectory representation for prediction purposes [43]. Some studies have proposed models for tasks ranging from predicting the next link using historical trajectories [44] to enhancing route prediction through pre-training and contrastive learning [21]. Furthermore, some models are designed for road network-constrained trajectory recovery, capable of recovering fine-grained points from low-sampling records [45, 46].
@@
II-B Knowledge Graph
II-B1 Knowledge Graph Completion
The rapidly expanding interest in KGs has fueled the advancement in tasks like recommender systems, question answering, and semantic search, given their ability to provide structured and machine-interpretable knowledge about real-world entities and their relations [47, 48, 49]. Despite their immense potential, a critical problem is the inherent incompleteness of information, making KGC an important and burgeoning research area. KGC refers to inferring missing or incomplete information in a KG by predicting new relationships between entities based on existing information [25].
Earlier studies on KGC typically employed statistical relational learning (SRL) methods, such as Markov Logic Networks (MLN) [50] and Probabilistic Soft Logic (PSL) [51]. These methods demonstrate effectiveness in capturing complex dependencies but need to improve scalability due to the need to specify all possible rules manually. More scalable machine learning approaches, especially those involving embeddings, have been proposed to overcome these limitations in recent years. TransE is a seminal model in this line, which models relations as translations in the entity embedding space [52]. Follow-up models such as TransH [23], TransR [53], and TransD [54] were subsequently proposed to handle complex relational data by introducing hyperplanes, relation-specific spaces, or dynamic mapping matrices respectively. Meanwhile, tensor factorization-based models like RESCAL [55], DistMult [56], and ComplEx [57] have been developed, aiming to capture the complex correlations between entities and relations. These models generally perform well but can be computationally intensive. More recently, models based on GNNs have shown promising results for KGC. Models such as R-GCN [58] and CompGCN [59] have achieved competitive results by modeling KGs as multi-relational graphs and learning from both the graph structure and node attributes.
@@ II-B1 Knowledge Graph Completion
II-B2 Mobility Knowledge Graph
@@
[60] devised a KG for urban traffic systems to uncover the implicit relationships amongst traffic entities and thereby unearth valuable traffic knowledge. Similarly, [61] constructed an urban movement KG using GPS trajectory data and affirmed the practicality of their model by predicting the level of user attention directed towards various city locations. [62] put forth a generalized framework for multi-source spatiotemporal data analysis, underpinned by KG embedding, intending to discern the network structure and semantic relationships embedded within multi-source spatiotemporal data. Several studies have focused on building KGs grounded on geographical information and human mobilities for various applications, such as predicting subsequent locations (i.e., Point of Interest recommendation) [63, 64, 65], modeling event streams [66], learning user similarity [67], forecasting destinations [68, 69], and performing epidemic contact tracing [70].
@@
Figure 1: The workflow of RouteKG.
III Preliminaries
In this section, we introduce definitions and the problem formulation in Section III-A.
III-A Problem Formulation
@@
A map-matched route $x$ of length $m$ is a sequence of links, $x=\left\{e^{1},e^{2},...,e^{m}\right\},x\in\mathcal{X}$. For every consecutive pair of links $(e^{i},e^{i+1})$, there exists a node $v$ in the graph $\mathbf{G}$ that connects the two edges. The set of all map-matched routes is denoted as $\mathcal{X}$. The $i$-th route can be partitioned into an observed route $x^{o}_{i}=\{e^{j}_{i}\}^{\Gamma}_{j=1}$ with length $\Gamma$ and a future route $x^{f}_{i}=\{e^{j}_{i}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$ with length $\Gamma^{\prime}$, where $\mathcal{X}^{o}=\{x_{i}^{o}\}_{i=1}^{|\mathcal{X}|}$ and $\mathcal{X}^{f}=\{x_{i}^{f}\}_{i=1}^{|\mathcal{X}|}$.
Given the above definitions, the short-term route prediction (or route prediction for short) problem can be broadly defined as the task of predicting the future route based on observed routes. However, as discussed in Section I, the availability of goal information plays a pivotal role in routing tasks. In some scenarios, no goal information is available. In other cases, we may know the rough direction of the destination, or its exact location. The degree of goal information inclusion can greatly influence the specific formulation of route prediction [71]. Remarkably, no existing work has undertaken an exhaustive evaluation encompassing all these distinct scenarios. Consequently, in this study, we categorize the route prediction problem into three subproblems: